# www.entalus.com

### v1

@@ 4.1 Refinement
Definition 4 (Stream Predicates)
A set $P\in\mathcal{P}({{A}^{\omega,{n}}})$ is also said to be a stream predicate. We simply write ${{{A}^{\omega,{n}}}}\!\!\restriction_{P}$ instead of the set comprehension $\{{f\in{{A}^{\omega,{n}}}}\,\mid\,{f\in P}\}$.
Definition 5 (Contracts)
Let $P\in{A}^{\omega}$ be a stream predicate, and $Q\,\coloneqq\,(Q_{f})_{f\in{A}^{\omega}}$ be an ${A}^{\omega}$-indexed family of stream predicate. Then, a contractual stream transformer $T$ has the (dependent) signature
$T$ $:(\Pi f\in{{{A}^{\omega,{n}}}}\!\!\restriction_{P})\,{{{B}^{\omega,{m}}}}\!\!\restriction_{Q_{f}}$ (17)
Now, $P$ is said to be the precondition of $T$ and $Q$ is the postcondition of $T$. The interface $[{P},{Q}]$ denotes the set of stream transformers with precondition $P$ and postcondition $Q$.
Whenever we write $T(f)$ we implicitly assume that this application is well-defined in that $f\in P$ and also that $T(f)\in Q_{f}$. Moreover, $[{P^{\prime}},{Q^{\prime}}]\subseteq[{P},{Q}]$ whenever $P\subseteq P^{\prime}$ and $Q^{\prime}\subseteq Q$ (pointwise). Notice also that contracts usually include fairness conditions for constraining nondeterminism [41].
5 Causality
We restrict our investigations to stream transformers whose outputs are determined by the history of inputs. The prefix ${f}\!\!\restriction_{n}$ of length $n\in\mathbb{N}$ for a stream $f$ is the finite word $([X^{0}]f)\ldots([X^{n-1}]f)$ of the first $n$ coefficients of $f$. Now, for a set $F$ of streams, ${F}\!\!\restriction_{n}$ denotes the set of prefixes ${f}\!\!\restriction_{n}$ for $f\in F$. Similarly, $[X^{k}]F$ denotes the set $\{{[X^{k}]f}\,\mid\,{f\in F}\}$. Causal stream transformers, at least in the single-valued case, are discussed in [11].
Definition 6 (Causal Stream Transformers)
The stream transformer $T:{{A}^{\omega,{n}}}\to\mathcal{P}({{B}^{\omega,{m}}})$ is $\delta$-causal, for $\delta\in\mathbb{N}$, if for all $k\in\mathbb{N}$ and $f,g\in{{A}^{\omega,{n}}}$
${f}\!\!\restriction_{k}={g}\!\!\restriction_{k}$  implies [$\text{~{}implies~{}}[X^{k+\delta}]T(f)=[X^{k+\delta}]T(g)\mbox{\@.}$ (18)
@@ 5 Causality
Every $\delta$-causal transformer is also $\delta^{\prime}$-causal for $\delta^{\prime}<\delta$. In particular, every strongly causal transformer is also weakly causal. This notion of strong causality is also closely related to the guardedness condition of de Bakker [4].
An alternative characterization of $\delta$-causality to the one in Definition 6 is easily established by natural induction on $n$.
Proposition 7
@@
${f}\!\!\restriction_{k}={g}\!\!\restriction_{k}$  $\text{~{}~{}implies~{}~{}}{T(f)}\!\!\restriction_{k+\delta}={T(g)}\!\!\restriction_{k+\delta}\mbox{\@.}$ (19)
Remark 12
Strongly causal stream transformer are often used for specifying clocked, hardware-like finite structures. In system design, however, working with the larger set of causal stream transformers is often preferred [6].
Example 15
A clocked $N$-bit register holding a value belonging to the alphabet $A\,\coloneqq\,\mathbb{B}^{N}$ may be viewed as a causal stream transformer ${(A\times\mathbb{B})}^{\omega}\to{A}^{\omega}$, where the second input $\mathbb{B}\,\coloneqq\,\{0,1\}$ corresponds to an enabling stream stating whether the incoming value on the first component ought to be loaded into the register or be ignored at a given time (or clock tick).
Example 16
The stream transformers from Example 16 are all (00-) causal, and $D_{1}$ is also $1$-causal.
Example 17 (Consing)
Consider, for $a\in A$, the family $Cons_{a}:{A}^{\omega}\to{A}^{\omega}$ of stream transformers, which map $z\mapsto a+X\cdot z$. These are $1$-causal transformers. $Cons_{0}$, in particular, yields the unit delay $D_{1}$ from Example 16.
If one is to implement a causal map by means of a machine that consumes successive values in an input stream, and produces a stream of successive values, it seems necessary that the map is continuous. Otherwise, the whole input stream would be needed at once, and an output would be forthcoming only at the end of time. Continuity means that finite information concerning the output of the function is determined by finite information concerning its input.
Example 18
The stream transformer $\mathrm{pos}(f)$ which returns the stream $1$ if all elements of $f$ are positive, and the stream 00 otherwise, is not causal.
5.1 Type 2 Computability
Type 2 Turing machines are extensions of Turing machine with a read-only input tape, a write-once output tape, and a working tape. Programs on a Type 2 machine run forever, generating increasingly more digits of the output [61]. The map $T$ is considered to be computable on a Type 2 machine if there is a total program, which takes finite time to write any number of symbols on the output tape regardless of the input. The working of a Type 2 machine up to any point of computation clearly only depends on a finite prefix of the input tape, and, therefore, every input tape that agrees with the given one up to that point would have caused the machine to write the same answer to the same output cells. Therefore all Type 2 computable functions are causal. As we will see below, causal functions are continuous. This implies the well-known fact that all Type 2 computable functions are continuous (see also Appendix 0.E).
5.2 Mealy Realizability
A Mealy machine is an intensional description of a causal stream transformer $T$, where the state holds enough information to determine $[X^{k}]T(f)$ from $[X^{k}]f$. More precisely, the set of weakly causal stream transformers ${A}^{\omega}\to{B}^{\omega}$ is isomorphic to the functions $A^{+}\to B$ from finite words over $A$ in $A^{+}$ to $B$, and a Mealy machine can be constructed for realizing such a word level function.
Example 19
[…]
$\delta((a,b),q)$$\,\coloneqq\,\begin{cases}(q,q)&\text{if }b=0\\ (a,a)&\text{if }b=1\mbox{\@.}\end{cases}$
Not every causal stream transformer, however, can be realized by a finite-state Mealy machine.
Example 20
Let $\mathrm{Prime}:{\mathbb{U}}^{\omega}\to{\mathbb{B}}^{\omega}$, where $\mathbb{U}$ is a singular set, be the Boolean-valued stream such that $[X^{k}]\mathrm{Prime}$ holds if and only if $k$ is a prime number. Now, the minimal Mealy machine realizing $\mathrm{Prime}$ has a countably infinite number of states.
5.3 Composition
Proposition 8
@@ 5.3 Composition
As a consequence, $\delta$-causal stream transformers, for $\delta>0$, are not preserved under composition, which results in an unfortunate inflexibility in system design. For instance, if we want to represent a stream transformer $T$ as a sequential composition $T_{1};T_{2}$ of two $1$-causal stream transformers then we always have to accept a delay by at least two. Both weakly and strongly causal stream transformers, however, are preserved, as a consequence of Proposition 8 under sequential and parallel composition.
Lemma 4
@@
5.4 Fixpoints
Strongly causal stream transformers have a unique fixpoint.
Example 21
For a $1$-causal stream transformer $T:{A}^{\omega}\to{A}^{\omega}$ one obtains a unique fixpoint $f^{*}\in{A}^{\omega}$ with $f^{*}=T(f^{*})$ by, recursively constructing Mealy machines whose transitions may also depend on outputs [46].
In contrast, fixpoints for weakly causal transformers may not exist and they may not be unique.
Example 22
  1. 1.
@@ 5.4 Fixpoints
Remark 13
Synchronous languages such as Lustre [25] or Esterel [7] therefore use syntactic restrictions for avoiding causal loops. The approach taken in the symbolic analysis laboratory SAL [5] relies on verification conditions asserting the absence of causal loops. The approach of SAL of semantically characterizing causality errors is more general. But it is also undecidable in general, since it can depend on arbitrary data properties.
Example 23
Let $c$, $d$, $e$ be Boolean-valued streams. Then a solution $(f,g)\in{(\mathbb{B}\times\mathbb{B})}^{\omega}$ of the stream equalities
@@
$T(f,g)$ $\,\coloneqq\,\mathrm{{c}\,?\,{(\lnot{g},e)}\,:\,{(d,f)}}$
There is no causal loop, because $f$ is causally dependent on $g$ only when $c$ is true, and vice-versa only when it is false. Therefore, these stream equations are acceptable in SAL.
{tikzpicture}
Figure 3: Sequential Composition.
{tikzpicture}
Figure 4: Parallel Composition with Copying and Merging.
6 Contraction
@@
Definition 7 (Contractions for Multivalued Maps)
@@
$d_{M}(u,v)$ $\leq l\cdot d_{N}(x,y)\mbox{\@.}$ (20)
If $l<1$ then the multivalued $T$ is said to be contractive, and if $l=1$ then $T$ is nonexpansive.
Remark 14
@@
From Lemma 5 one obtains metric-based characterizations for weak and also for strong causality.
Corollary 3
A stream transformer $T:{{A}^{\omega,{n}}}\to\mathcal{P}({{B}^{\omega,{m}}})\setminus\emptyset$ is
  1. 1.
    weakly causal if and only if it is nonexpansive, and
  2. 2.
@@
6.1 Fixpoints of Stream Transformers
Recall that a strictly contracting or shrinking map $T$ satisfies $d(T(f),T(g))<d(f,g)$ for all $f,g$. Now, any shrinking map in a spherically complete ultrametric space has a unique fixpoint ([43], Theorem 1).
Definition 8
For $T:{{A}^{\omega,{n}}}\to{{A}^{\omega,{n}}}$ a strictly contractive stream transformer, ${T}^{+}$ is the unique fixpoint of ${\cal T}\,\coloneqq\,(\lambda S)\,S;T$.
${T}^{+}$ is well-defined by ([43], Theorem 1), since, using Lemma 2, maps of (deterministic) stream transformers form an ultrametric and spherically complete space, and ${\cal T}$ is strictly contractive in this space.
$d({\cal T}(S_{1}),{\cal T}(S_{2}))$ $=d((S_{1};T),(S_{2};T))$
@@ 6.1 Fixpoints of Stream Transformers
Using fixpoint results for strictly contractive multivalued maps (Theorem 2 in [35]) Definition 8 is generalized to obtain a unique fixpoint of transformers of strictly contracting multivalued stream transformers.
6.2 Lipschitz Contraction
We show that contraction on multivalued maps with nonempty compact codomains coincides with contraction with respect to the Hausdorff metric as defined in [38].
Definition 9 (Lipschitz Contraction)
@@ 6.2 Lipschitz Contraction
$\mathcal{H}_{d_{N}}({T(x)},{T(y)})\leq l\cdot d_{M}(x,y)\mbox{\@,}$ (21)
@@
from $d(x,y)<\delta$.
The continuous image of a compact set is compact. Assume an open cover of $T(K)$. As $T$ is continuous, the inverse image of those open sets form an open cover for $K$. Since $K$ is compact, there is a finite subcover of $T(K)$, and, by construction, the images of the finite subcover give a finite subcover of $T(K)$, and therefore $T(K)$ is compact.
Lemma 6
@@
We present some elementary results which will be used in later sections. The proofs of many of these facts are straightforward (see also Proposition 8).
Proposition 9
@@
Proposition 10 gives a technique for constructing a Lipschitz mapping from a finite number of single-valued Lipschitz mapping by ”unioning their graphs at each point”. The closure condition (10) of Lipschitz maps under angelic nondeterminism $S\mathrm{\sqcup}T$ can be generalized to an arbitrary ${\cal I}$-index family $(T_{\iota})_{\iota\in{\cal I}}$.
Remark 16
The set of Lipschitz maps, however, is not closed under demonic nondeterminism $(S\mathrm{\sqcap}T)$, which maps $x\mapsto S(x)\,\cap\,T(x)$ (assuming all these intersections are non-empty). Indeed, the demonic combination of Lipschitz maps might not even be continuous [38].
For stream transformers with non-empty compact images, Lipschitz contraction (Definition 9) is equivalent to the notion of contraction mappings in Definition 7.
Lemma 7
@@
    By compactness of $T(y)$,111 For every compact topological space $M$ and $f:M\to\mathbb{R}$ continuous: the image $f(M)$ is bounded and there exists $a,b\in M$ such that $f(a)=\inf_{x\in M}f(x)$ and $f(b)=\sup_{x\in M}f(x)$. there exists $v_{0}\in T(y)$ such that
@@
Remark 17
The proof of the left-to-right statement of Lemma 7 only requires the non-empty images of $T$ to be closed (and bounded), and not necessarily compact.
Lemma 7 and Corollary 3 together yield the correspondence between causal and Lipschitz contractive maps.
Corollary 4
The stream transformer $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{B}^{\omega,{m}}}})}$ is
  1. 1.
    weakly causal if and only if it is Lipschitz nonexpansive, and
  2. 2.
@@
7 Fixpoints
We have already seen how fixpoints for strictly causal functions can be constructed using recursive Mealy machines (Example 21). We pursue, however, a less syntax- and machine-oriented path, for solving stream equations, and we compute fixpoints for vector- and multivalued maps.
@@ 7.1 Set Transformers
    ${\mathrm{sp}}_{T}(P)$ $\,\coloneqq\,T(P)\mbox{\@;}$ (23)
  2. 2.
@@
    ${\mathrm{wp}}_{T}(Q)$$\,\coloneqq\,\{{x\in{\mathrm{Comp}({M})}}\,\mid\,{T(x)\subseteq Q}\}\mbox{\@.}$ (24)
It follows that the ${\mathrm{sp}}_{T}$ and ${\mathrm{wp}}_{T}$ are adjoint.
${\mathrm{sp}}_{T}(P)\subseteq Q$ $\,\iff\,P\subseteq{\mathrm{wp}}_{T}(Q)$ (25)
@@
Lemma 8 (WP Stream Calculus)
@@
Remark 18
A Hoare-like stream contract $\{P\}\,T\,\{Q\}$ for a nondeterministic stream transformer $T$ with precondition $P$ and postcondition $Q$ holds if and only if $P\subseteq{\mathrm{wp}}_{T}(Q)$, or, equivalently, ${\mathrm{sp}}_{T}(P)\subseteq Q$. Corresponding rules of a Hoare-like calculus are derived from the weakest precondition calculus (Lemma 8).
Proposition 11
@@
Remark 19
Compared with the usual Knaster-Tarski fixpoint iteration of Scott-continuous transformers, the main advantage of the iteration in Lemma 9 is that, at any iteration, there is a quantitative measure of the distance to the fixpoint and also of the progress towards this fixpoint. On the other hand, the iterations in Lemma 9 generally neither are under- nor overapproximations of fixpoints.
Lemma 10
@@
Proof
For a fixpoint $f^{*}$ of $T$, that is, $f^{*}\in T(f^{*})$ define the iteration $F_{0}\,\coloneqq\,\{f^{*}\}$ $F_{k+1}\,\coloneqq\,{\mathrm{sp}}_{T}(F_{k})=\{{g\in T(f)}\,\mid\,{f\in F_{k}}\}$. By induction on $k$, $f^{*}\in F_{k}$ for each $k\in\mathbb{N}$, and therefore $f^{*}\in\lim_{k\to\infty}F_{k}={\mathrm{fix}}({{\mathrm{sp}}_{T}})$.
The subset relation in Lemma 10 may be strict [38].
Example 24 ([38])
For simplicity, we work in the unit interval $[0,1]$ with the usual Euclidean metric. Let
$t(x)\,\coloneqq\,\begin{cases}\nicefrac{{1}}{{2}}(x+1):\text{if }0\leq x\leq\nicefrac{{1}}{{2}}\\ (-\nicefrac{{1}}{{2}})x+1:\text{if }\nicefrac{{1}}{{2}}<x\leq 1\end{cases}$
and the multivalued $T:I\to{\mathrm{Comp}({I})}$ maps $x\mapsto\{0\}\,\cup\,\{t(x)\}$. Evidently, $T$ is a multivalued contraction, the fixpoints of the multivalued $T$ are $\{0,\nicefrac{{2}}{{3}}\}$, and the fixpoint of the set transformer ${\mathrm{sp}}_{T}$ induced by $T$ is $\{0,\nicefrac{{2}}{{3}},t(0),t^{2}(0),t^{3}(0),\cdots\}$.
When applying Lemma 9 to streams ${{A}^{\omega,{n}}}$ and its ultrametric $d$, then the convergence bounds can be improved to $l^{k}$ with $l=(\nicefrac{{1}}{{2}})^{\delta}$ for some $\delta>0$, since $d$ is bounded and the application of the strengthened triangle inequality in the proof of Banach’s contraction principle yields the improved bound $l^{k}$.
Theorem 7.1
Let $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{B}^{\omega,{m}}}})}$ be $\delta$-causal, for $\delta>0$. With the notation and the iterations $F_{k}$, $G_{k}$ as in Lemma 9, the unique fixpoints ${\mathrm{fix}}({{\mathrm{sp}}_{T}})$ and ${\mathrm{fix}}({{\mathrm{wp}}_{T}})$ are obtained as the limits of $F_{k}$ and $G_{k}$, respectively, as $k\to\infty$. Furthermore, for $k\in\mathbb{N}$:

### v2

@@ 4. Transformers
5. Causality
We restrict our investigation to stream transformers whose outputs are determined by their history of inputs. The prefix ${f}\!\!\restriction_{n}$ of length $n\in\mathbb{N}$ for a stream $f$ is the finite word $([X^{0}]f)\ldots([X^{n-1}]f)$ of the first $n$ coefficients of $f$. Now, for a set $F$ of streams, ${F}\!\!\restriction_{n}$ denotes the set of prefixes ${f}\!\!\restriction_{n}$ for $f\in F$. Similarly, $[X^{k}]F$ denotes the set $\{{[X^{k}]f}\,\mid\,{f\in F}\}$. Causal stream transformers, at least in the single-valued case, are discussed in (Broy and Stølen, 2012).
Definition 5.1 (Causal Stream Transformers).
The nondeterministic stream transformer $T:{{A}^{\omega,{n}}}\to\mathcal{P}({{B}^{\omega,{m}}})$ is $\delta$-causal, for $\delta\in\mathbb{N}$, if for all $k\in\mathbb{N}$ and $f,g\in{{A}^{\omega,{n}}}$
(16) ${f}\!\!\restriction_{k}={g}\!\!\restriction_{k}$  implies [$\text{~{}implies~{}}[X^{k+\delta}]T(f)=[X^{k+\delta}]T(g)\mbox{\@.}$
@@ 5. Causality
Every $\delta$-causal transformer is also $\delta^{\prime}$-causal for $\delta^{\prime}<\delta$. In particular, every strongly causal transformer is also weakly causal.
An alternative characterization of $\delta$-causality to the one in Definition 5.1 is easily established by natural induction on $k$.
Proposition 5.2.
@@
(17) ${f}\!\!\restriction_{k}={g}\!\!\restriction_{k}$  $\text{~{}~{}implies~{}~{}}{T(f)}\!\!\restriction_{k+\delta}={T(g)}\!\!\restriction_{k+\delta}\mbox{\@.}$
Strongly causal stream transformer are often used to specify clocked, hardware-like finite structures. However, when designing a system, it is often preferable to work with the larger class of weakly causal stream transformers (Benveniste and Berry, 1991).
Example 5.3.
A clocked $N$-bit register holding a value belonging to the alphabet $A\,\coloneqq\,\mathbb{B}^{N}$ induces a causal stream transformer ${(A\times\mathbb{B})}^{\omega}\to{A}^{\omega}$, where the second input $\mathbb{B}\,\coloneqq\,\{0,1\}$ corresponds to an enabling stream for stating whether the incoming value on the first component ought to be loaded into the register or be ignored at a given clock tick.
Example 5.4.
The basic stream transformers in Example 4.2 for constructing stream circuits are all (00-) causal, and $D_{1}$ is also $1$-causal.
Example 5.5 (Consing).
Consider, for $a\in A$, the family $Cons_{a}:{A}^{\omega}\to{A}^{\omega}$ of stream transformers, which map $z\mapsto a+X\cdot z$. These are $1$-causal transformers. $Cons_{0}$, in particular, yields the unit delay $D_{1}$ from Example 5.4.
Example 5.6.
The stream transformer which returns the stream $1$ if all elements of its argument stream are positive, and the stream 00 otherwise, is not causal.
Mealy Realizability.
A Mealy machine is an intensional description of a causal stream transformer $T$, where the state holds enough information to determine $[X^{k}]T(f)$ from $[X^{k}]f$. More precisely, the set of weakly causal stream transformers ${A}^{\omega}\to{B}^{\omega}$ is isomorphic to the functions $A^{+}\to B$ from finite words over $A$ in $A^{+}$ to $B$, and a Mealy machine, possibly containing an infinite number of states, can be constructed for realizing such a word level function.
Composition.
Causality is compositional for the basic operators on transformer from Definition 4.5.
Proposition 5.7.
@@
As a consequence, $\delta$-causal stream transformers, for $\delta>0$, are not preserved under composition, which results in an unfortunate inflexibility in system design. For instance, if we want to represent a stream transformer $T$ as a sequential composition $T_{1};T_{2}$ of two $1$-causal stream transformers then we always have to accept a delay by at least two. Both weakly and strongly causal stream transformers, however, are preserved, as a consequence of Proposition 5.7 under sequential and parallel composition.
Lemma 5.8.
@@
Fixpoints.
Strongly causal stream transformers have a unique fixpoint. This can be shown, for example, by constructing recursive Mealy machines whose transitions may also depend on outputs (Pradic, 2020). In contrast, fixpoints for weakly causal transformers may not exist and they may not be unique.
Example 5.9.
  1. (1)
@@
Synchronous languages such as Lustre (Halbwachs, 2012) or Esterel (Berry and Gonthier, 1992) therefore use syntactic restrictions for avoiding causal loops. In contrast, the approach taken in the symbolic analysis laboratory SAL (Bensalem et al., 2000) relies on verification conditions asserting the absence of causal loops.
Example 5.10.
Let $c$, $d$, $e$ be Boolean-valued streams. Then a solution $(f,g)\in{(\mathbb{B}\times\mathbb{B})}^{\omega}$ of the stream equalities111 $\mathrm{{c}\,?\,{f}\,:\,{g}}$ denotes the (pointwise) conditional stream ”if $c$ then $f$ else $g$”.
@@
$T(f,g)$ $\,\coloneqq\,\mathrm{{c}\,?\,{(\lnot{g},e)}\,:\,{(d,f)}}\mbox{\@.}$
There is no causal loop, because $f$ is causally dependent on $g$ only when $c$ is true, and vice versa only when it is false. Therefore, these stream equations are acceptable in SAL.
The approach of SAL of semantically characterizing causality errors is more general than the synctactic restrictions, say, in Lustre. But it is also undecidable in general, since it can depend on arbitrary data properties.
6. Contraction
@@
6.1. Multivalued Contraction
Definition 6.1 (Contractions for Multivalued Maps).
@@ 6.1. Multivalued Contraction
(18) $d_{M}(u,v)$ $\leq l\cdot d_{N}(x,y)\mbox{\@.}$
If $l<1$ then the multivalued $T$ is said to be (strictly) contractive, and if $l=1$ then $T$ is nonexpansive.
@@
This finishes the proof. ∎
From Lemma 6.2 one obtains metric-based characterizations for weak and also for strong causality.
Corollary 6.3 (Causality-Contraction Connection).
A nondeterministic stream transformer $T:{{A}^{\omega,{n}}}\to\mathcal{P}({{B}^{\omega,{m}}})\setminus\emptyset$ is:
  1. (1)
    Weakly causal if and only if it is nonexpansive.
  2. (2)
@@
Note that this correspondence can be extended for shrinking transformers together with a new notion of causality with non-uniform bounds on determined outputs.
6.2. Fixpoints of Stream Transformers
Recall that a strictly contracting or shrinking map $T$ satisfies the strict inequality $d(T(f),T(g))<d(f,g)$ for all $f,g$. Now, any shrinking map in a spherically complete ultrametric space has a unique fixpoint ((Petalas and Vidalis, 1993), Theorem 1).
Definition 6.4.
For $T:{{A}^{\omega,{n}}}\to{{A}^{\omega,{n}}}$ a strictly contractive stream transformer, ${T}^{+}$ is the unique fixpoint of ${\mathcal{T}}\,\coloneqq\,(\lambda S)\,S;T$.
${T}^{+}$ is well-defined by ((Petalas and Vidalis, 1993), Theorem 1), since, using Lemma 3.3, maps of (deterministic) stream transformers form an ultrametric and spherically complete space, and ${\mathcal{T}}$ is strictly contractive in this space.
$d({\mathcal{T}}(S_{1}),{\mathcal{T}}(S_{2}))$ $=d((S_{1};T),(S_{2};T))$
@@ 6.2. Fixpoints of Stream Transformers
Using fixpoint results for strictly contractive multivalued maps (Theorem 2 in (Kubiaczyk and Ali, 1996)) Definition 6.4 is generalized to obtain a unique fixpoint of transformers of strictly contracting multivalued stream transformers.
6.3. Lipschitz Contraction
We show that contraction on multivalued maps with nonempty compact codomains coincides with contraction with respect to the Hausdorff metric as defined in (Nadler, 1969).
Definition 6.5 (Lipschitz Contraction).
@@ 6.3. Lipschitz Contraction
(19) $\mathcal{H}_{d_{N}}({T(x)},{T(y)})\leq l\cdot d_{M}(x,y)\mbox{\@,}$
@@
from $d(x,y)<\delta$. The continuous image of a compact set is compact. Assume an open cover of $T(K)$. As $T$ is continuous, the inverse image of those open sets form an open cover for $K$. Since $K$ is compact, there is a finite subcover of $T(K)$, and, by construction, the images of the finite subcover give a finite subcover of $T(K)$, and therefore $T(K)$ is compact.
Lemma 6.6.
@@
We present some elementary results which will be used in later sections ((Nadler, 1969)). The proofs of many of these facts are straightforward (see also Proposition 5.7).
Proposition 6.7.
@@
Proposition 6.8 gives a technique for constructing a Lipschitz mapping from a finite number of single-valued Lipschitz mapping by ”unioning their graphs at each point”. The closure condition (6.8) of Lipschitz maps under angelic nondeterminism $S\mathrm{\sqcup}T$ can be generalized to an arbitrary ${\mathcal{I}}$-index family $(T_{\iota})_{\iota\in{\mathcal{I}}}$ (Nadler, 1969).
For stream transformers with nonempty compact images, Lipschitz contraction (Definition 6.5) is equivalent to the notion of contraction mappings in Definition 6.1.
Lemma 6.9.
@@
    By compactness of $T(y)$, there exists $v_{0}\in T(y)$ such that
@@
The proof of the left-to-right statement of Lemma 6.9 only requires the nonempty images of $T$ to be closed (and bounded), and not necessarily compact. Lemma 6.9 and Corollary 6.3 together yield the correspondence between causal and Lipschitz contractive maps.
Corollary 6.10 (Causality-Contraction Connection (Lipschitz)).
The nondeterministic stream transformer $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{B}^{\omega,{m}}}})}$ is:
  1. (1)
    Weakly causal if and only if it is Lipschitz nonexpansive.
  2. (2)
@@
7. Fixpoints
Fixpoints for strictly causal functions are constructed, for example, on the basis of recursive Mealy machines (Pradic, 2020). We pursue, however, a less syntax- and machine-oriented path, for solving stream equations, and we compute fixpoints for vector- and multivalued maps.
@@ 7.1. Set Transformers
    (21) ${\mathrm{sp}}_{T}(P)$ $\,\coloneqq\,T(P)\mbox{\@;}$
  2. (2)
@@
    (22) ${\mathrm{wp}}_{T}(Q)$$\,\coloneqq\,\{{x\in{\mathrm{Comp}({M})}}\,\mid\,{T(x)\subseteq Q}\}\mbox{\@.}$
It follows that the ${\mathrm{sp}}_{T}$ and ${\mathrm{wp}}_{T}$ are adjoint in that
(23) ${\mathrm{sp}}_{T}(P)\subseteq Q$ $\,\iff\,P\subseteq{\mathrm{wp}}_{T}(Q)$
@@
Example 7.2 (WP Stream Calculus).
@@
Example 7.3.
A Hoare-like stream contract $\{P\}\,T\,\{Q\}$ for a nondeterministic stream transformer $T$ with precondition $P$ and postcondition $Q$ holds if and only if $P\subseteq{\mathrm{wp}}_{T}(Q)$, or, equivalently, ${\mathrm{sp}}_{T}(P)\subseteq Q$. Corresponding rules for a Hoare-like calculus are derived from the weakest precondition calculus (Lemma 7.2).
Proposition 7.4.
@@
Compared with the usual Knaster-Tarski fixpoint iteration of Scott-continuous transformers, the main advantage of the iteration in Lemma 7.5 is that, at any iteration, there is a quantitative measure of the distance to the fixpoint and also of the progress towards this fixpoint. On the other hand, the iterations in Lemma 7.5 generally neither are under- nor overapproximations of fixpoints.
Lemma 7.6.
@@
Proof.
For a fixpoint $f^{*}$ of $T$, that is, $f^{*}\in T(f^{*})$ define the iteration $F_{0}\,\coloneqq\,\{f^{*}\}$ $F_{k+1}\,\coloneqq\,{\mathrm{sp}}_{T}(F_{k})=\{{g\in T(f)}\,\mid\,{f\in F_{k}}\}$. By induction on $k$, $f^{*}\in F_{k}$ for each $k\in\mathbb{N}$, and therefore $f^{*}\in\lim_{k\to\infty}F_{k}={\mathrm{fix}}({{\mathrm{sp}}_{T}})$. ∎
The subset relation in Lemma 7.6 may be proper (Nadler, 1969). When applying Lemma 7.5 to streams ${{A}^{\omega,{n}}}$ and its ultrametric $d$, then the convergence bounds can be improved to $l^{k}$ with $l=(\nicefrac{{1}}{{2}})^{\delta}$ for some $\delta>0$, since $d$ is bounded and the application of the strengthened triangle inequality in the proof of Banach’s contraction principle yields the improved bound $l^{k}$.
Theorem 7.7.
Let $T:{{A}^{\omega,{n}}}\to{\mathrm{Comp}({{{B}^{\omega,{m}}}})}$ be $\delta$-causal, for $\delta>0$. With the notation and the iterations $F_{k}$, $G_{k}$ as in Lemma 7.5, the unique fixpoints ${\mathrm{fix}}({{\mathrm{sp}}_{T}})$ and ${\mathrm{fix}}({{\mathrm{wp}}_{T}})$ are obtained as the limits of $F_{k}$ and $G_{k}$, respectively, as $k\to\infty$. Furthermore, for $k\in\mathbb{N}$: